OpenStudy (freemap):

Please Check Me Answer, I'm Not Completely Sure Which represents the first three terms of the sequence: a1 = 2 and an = 4(an-1)2? A. 2, 16, 1024 B. 4, 16, 1024 C. 2, 16, 32 D. 4, 16, 32 My answer is A?

jimthompson5910 (jim_thompson5910):

Correct first term = 2 second term = 2*(first term)^2 = 2*2^2 = 16 third term = 2*(second term)^2 = 2*16^2 = 1024

OpenStudy (freemap):

Thanks so much

jimthompson5910 (jim_thompson5910):

no problem